OpenStudy (anonymous):

integral of 1/(sqrt(x)+cube root of x)

OpenStudy (anonymous):

\[\int \frac{\text{d}x}{\sqrt{x}+\sqrt[3]{x}}\]right?

OpenStudy (anonymous):

yes. Sorry..I didnt know how to write it another way

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

have u tried it yet?

OpenStudy (anonymous):

the problem says to make a substitution to express the integrand as a rational function and then avaluate the integral.

OpenStudy (anonymous):

I have, but am not sure how to go about with the substitution

OpenStudy (anonymous):

using substitution \[x=t^6\]

OpenStudy (anonymous):

the book gave a hint and is saying to use u=\[\sqrt[6]{x?}\]

OpenStudy (anonymous):

but i dont know how they got that.

OpenStudy (anonymous):

or where to plug it in

OpenStudy (anonymous):

yes its same\[u=\sqrt[6]{x}\]\[u^6=x\]u use that sub to get rid of radicals

OpenStudy (loser66):

perfect instruction, hand off, @mukushla

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

so would it be 1/(u^6+u^1/2) ?

OpenStudy (anonymous):

thanks @Loser66 looking more carefully\[\int \frac{6t^5\text{d}t}{t^3+t^2}=\int \frac{6t^3\text{d}t}{t+1}\]

OpenStudy (anonymous):

sorry, its better to work with ur notation\[\int \frac{6u^5\text{d}u}{u^3+u^2}=\int \frac{6u^3\text{d}u}{u+1}\]

OpenStudy (anonymous):

if you have x=u^6, then wouldnt \[\sqrt[3]{x}\] with x^6 inside be x^1/2?

OpenStudy (anonymous):

u^6 will be inside not x^6 :)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

im not sure how you got the u^2 on the denominator

OpenStudy (anonymous):

well\[\sqrt[3]{x}=\sqrt[3]{u^6}=(u^6)^{\frac{1}{3}}=u^{\frac{6}{3}}=u^2\]

OpenStudy (anonymous):

makes sense? :)

OpenStudy (anonymous):

oh i see it now..thanks alot!

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

and you got the 6u^5 by taking the derivative of u right?

OpenStudy (loser66):

yes.

OpenStudy (loser66):

\[\int \frac{6u^3\text{d}u}{u+1}\] step up from this by long division and very easy to get the answer then, right?

OpenStudy (anonymous):

yes i did long division and then used partial fractions

OpenStudy (loser66):

what do you get from long division?

OpenStudy (anonymous):

u^2-u+1-(1/u+1)

OpenStudy (loser66):

yep, for the last one, no need to use partial, you have formula for this, =ln|u+1|

OpenStudy (loser66):

don't forget, replace u by x at the final answer

OpenStudy (anonymous):

i did. Thanks alot!

OpenStudy (loser66):

give medal to mukusla, he gave you a perfect instruction. beautiful substitution. !!!